The sum of two consecutive integers is one less than twice the greater integer. Find the integers.
Answer by DrBeeee(684)

Let n = the first number
Then the next consecutive number is (n+1).
Apply your problem statement and get
(1) n + (n+1) = 2*(n+1) -1 or
(2) 2*n + 1 = 2*n + 2 -1 or
(3) 2*n = 2*n or
(4) n = n
Since this is an identity ie n=n, your answer can be any two consecutive intergers. Try it.
Use 8,9.
Is (8+9 = 2*9-1)?
Is (17 = 17)? Yes
Use -99,-98
Is (-98 +(-99) = 2*(-98)-1)?
Is (-197 = -196-1)?
Is (-197 = -197)? Yes
Try any two and it works!!